# Optimal Online Change Detection via Random Fourier Features

### v1

@@ 5 Minimax optimality of RFF-MMD
Recall that with the conditions of Assumption 1, the underlying kernel is characteristic and $\operatorname{MMD}_{K}$ metricizes the space $\mathcal{M}_{1}^{+}$. Therefore, $\operatorname{MMD}_{K}[\mathbb{P},\mathbb{Q}]>0$ for any $\mathbb{P}\neq\mathbb{Q}$, and (3) guarantees that our stopping time (10) obtains a finite detection delay for any fixed alternative. Still, one may ask whether the detection delay is optimal. The following theorem resolves this question and shows that the detection delay of RFF-MMD is essentially optimal from a minimax perspective, up to logarithmic terms.
@@
For every continuous, bounded, and shift invariant kernel $K:\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}$ there is a constant $C_{K}$ depending only on $K$, and absolute constants $\alpha_{0},\beta_{0}\in(0,1)$ independent of $K$, such that for any $\alpha\leq\alpha_{0}$ it holds that
@@ 6 Experiments
We collect our experiments on synthetic data in Section 6.1 and on the MNIST data set in Section 6.2. We refer to the supplement for a numerical comparison of the different thresholds for the stopping rule. To interpret the change point detection performance of the proposed method, we compare its average run length (ARL) and expected detection delay (EDD) to the existing kernel-based methods presented in Table 1.333 All code replicating the experiments is available at https://github.com/FlopsKa/rff-change-detection. For all experiments, we use the Gaussian kernel $K(\mathbf{x},\mathbf{y})=e^{-\gamma\left\|{\mathbf{x}-\mathbf{y}}\right\|_{2}^{2}}$ with $\gamma$ set by the median heuristic [11] or its RFF approximation, depending on the algorithm. All results were obtained on a PC with Ubuntu 20.04 LTS, 124GB RAM, and 32 cores with 2GHz each.
@@ 6.1 Synthetic data
Runtime. Figure 1 summarizes the runtime results of Algorithm 1 with random Fourier features $r\in\{10,50,100,500,1{.}000\}$ and for streams of length up to $n=250{.}000$. The experiments verify the $\mathcal{O}(r\log n)$ runtime complexity of the proposed algorithm, derived analytically in Section 4.2. We note that the dependence on $d$ is linear; we consider $d=1$ only.
@@
[…]
@@ 6.2 MNIST data
In this section, we interpret the MNIST data set [8] as high-dimensional data stream, similar to Wei and Xie [55, Figure 7], with the goal of detecting a change when the digit changes from 00 to a different digit. The experimental setup matches that of Section 6.1, but with $d=784$. Figure 3 collects our results; the results for the digits 4–6 are similar and in the supplement. Similar to Figure 2, the proposed Online RFF-MMD algorithm shows very good performance. In particular, it achieves a lower EDD than all tested competitors throughout.
Figure 3: Average detection delay from MNIST digit 0 to digits $1$, $2$, and $3$ (left to right).
@@ 7 Limitations
As with all kernel-based tests, the choice of kernel impacts the power of the test. While our theoretical guarantees (Section 4.3) hold for any kernel satisfying Assumption 1 and do not require any pre-change data, one usually selects the kernel or its parameters using a few available samples in practice. While kernel optimization is not the focus of this work, there exist works [19, 20, 21, 31, 46, 47, 15, 16, 17] to (approximately) achieve this goal; it is interesting future work to tackle this problem in the sequential setting.
@@
The authors thank Zoltán Szabó for helpful discussions. FK thanks Georg Gntuni and Marius Bohnert for exchanges on the algorithm’s implementation. This work was supported by the pilot program Core-Informatics of the Helmholtz Association (HGF).
@@ Appendix A Appendix
This supplementary material is structured as follows. We detail the extension of Online RFF-MMD to take a known or observed pre-change distribution into account in Appendix A.1. In particular, we show that in these settings tighter thresholds are possible. In Appendix A.2, we provide a schematic representation of Algorithm 1. Additional numerical results of MMD-RFF are in Appendix A.3; we include numerical results w.r.t. our tighter thresholds in Appendix A.3.3. Appendix A.4 collects all our proofs.
A.1 Extensions: known or estimable pre-change distribution
@@ A.1.1 Incorporating information of the pre-change distribution
To begin with, we consider the setting in which, for some $\nu\in\mathbb{N}$, historical data $X_{-\nu+1},\dots,X_{0}$ known to be from the (nonetheless unknown) pre-change distribution $\mathbb{P}$ is available. This setting has been studied in the literature from both the parametric [40] and non-parametric [55] perspectives. It is straightforward to extend the Online RFF-MMD stopping time defined in the main text to take advantage of the additional information. Intuitively, for each local test one may prepend the historical data to the block of data take to be from the pre-change distribution. More formally, the following stopping time can be used:
$N=\inf\left\{n\geq 2\mid\bigcup_{j=1}^{\left\lfloor\log_{2}(n)\right\rfloor-1}\sqrt{\frac{2^{j}(n+\nu-2^{j})}{n+\nu}}\operatorname{MMD}_{\hat{K}}\!\left[X_{(-\nu+1):(n-2^{j})},X_{(n-2^{j}+1):n}\right]>\lambda_{n}\right\}.$
@@
$N=\inf\left\{n\geq 2\mid\bigcup_{j=1}^{\left\lfloor\log_{2}(n)\right\rfloor-1}\sqrt{2^{j}}\operatorname{MMD}_{\hat{K}}\!\left[\mathbb{P},X_{(n-2^{j}+1):n}\right]>\lambda_{n}\right\}$
@@ A.1.2 Sharper thresholds through knowledge of the pre-change distribution
The main tool for proving Theorems 1 and 2 is Lemma 5, which controls the tail behavior of the test performed by Online RFF-MMD under their respective local nulls. With additional knowledge of the pre-change distribution, Lemma 5 can be significantly sharpened by taking the second moment of the feature map into account. To that end, we have the following result.
@@
$\mathbb{P}\!\left(\operatorname{MMD}_{\hat{K}}\!\left[X_{1:n},Y_{1:m}\right]>\varepsilon\right)\leq 2\exp\left(-\frac{1}{2}\min(m,n)\varepsilon^{2}\left[\tilde{\sigma}^{2}+2\sqrt{2}\varepsilon\right]^{-1}\right),$
where $\tilde{\sigma}^{2}=2\mathbb{E}_{X\sim\mathbb{P}}\left[K\left(X,X\right)\right]-\mathbb{E}_{X,Y\sim\mathbb{P}}\left[K\left(X,Y\right)\right]$
@@
$\lambda_{n,j}=\frac{4\sqrt{2}f(\gamma)}{\sqrt{\min(2^{j},n-2^{j})}}+\tilde{\sigma}\sqrt{2f(\gamma)},\hskip 20.00003ptn\in\mathbb{N},j\leq\log_{2}(n)$ (14)
$f(\gamma)=\log(4\gamma\log_{2}(2\gamma))$
it holds that $\mathbb{E}_{\infty}[N]\leq\gamma$ where $N$ is as defined in (10).
@@
$\lambda_{n,j}=\frac{4\sqrt{2}f(\alpha,n)}{\sqrt{\min(2^{j},n-2^{j})}}+\tilde{\sigma}\sqrt{2f(\alpha,n)},\hskip 20.00003ptn\in\mathbb{N},j\leq\log_{2}(n)$
$f(\alpha,n)=\log(n/\alpha)+\log(\log_{2}(n))+\frac{1}{2}\log(\log_{2}(n))$
it holds that $\mathbb{P}_{\infty}(N<\infty)\leq\gamma$ where $N$ is as defined in (10).
@@
A.2 Further details on the Online RFF-MMD algorithm
Figure 4 illustrates Algorithm 1. The details are as follows. Upon observing the first element $X_{1}$, the algorithm creates a new Window $W$, storing the feature map $\hat{z}_{K}(X_{1})$ and that it has one element. Similarly, upon observing $X_{2}$, the algorithm creates a new window $W^{\prime}$, storing $\hat{z}_{K}(X_{2})$ and $c=1$. As both windows $W,W^{\prime}$ have the same counts, the algorithm merges them into a new window $W$, storing $\hat{z}_{K}(X_{1})+\hat{z}_{K}(X_{2})$ and $c=2$. The algorithm proceeds in this manner, resulting in the construction of the dyadic grid outlined in Section 4. For example, when observing $X_{6}$, there are again two windows of size $1$, which the algorithm merges to store a total of two windows, one capturing $X_{1},\ldots,X_{4}$ and the other one capturing $X_{5},X_{6}$. We recall that the observations themselves are never stored explicitly.
@@ A.3 Additional experiments
In this section, we summarize additional numerical results. The MNIST results with the MC threshold as in the main text are in Section A.3.1. We include results obtained without threshold estimation in Section A.3.2. A comparison of the distribution dependent and distribution-free bounds is in Section A.3.3.
A.3.1 MNIST digits 4–9
In the following Figure 5, we collect the change detection performances of the algorithms of Table 1 on MNIST data, where the digit changes from 00 to one of 44449999; the results on the other digits and the experimental setup are in Section 6.2. As in the corresponding experiment in the main part of this article, our proposed Online RFF-MMD algorithm consistently achieves the lowest expected detection delay, highlighting its good practical performance.
@@ A.3.1 MNIST digits 4–9
A.3.2 Distribution-free bound

### v2

@@ 5 Minimax optimality of RFF-MMD
Recall that with the conditions of Assumption 1, the underlying kernel is characteristic and $\operatorname{MMD}_{K}$ metrizes the space $\mathcal{M}_{1}^{+}$. Therefore, $\operatorname{MMD}_{K}[\mathbb{P},\mathbb{Q}]>0$ for any $\mathbb{P}\neq\mathbb{Q}$ and we have that (11) guarantees that our stopping time (8) obtains a finite detection delay, with high probability for any fixed alternative. Still, one may ask whether the detection delay is optimal. The following theorem resolves this question and shows that the detection delay of RFF-MMD is essentially optimal from a minimax perspective, up to logarithmic terms.
@@
For every kernel $K:\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}$ satisfying Assumption 1 there is a constant $C_{K}$ depending only on $K$ and absolute constants $\alpha_{0},\beta_{0}\in(0,1)$ independent of $K$, such that for any $\alpha\leq\alpha_{0}$ it holds that
@@ 6 Experiments
We collect our experiments on synthetic data in Section 6.1 and on the MNIST data set in Section 6.2. We refer to Appendix A.3 for additional experiments and a numerical comparison of different thresholds for the stopping rule. To interpret the change point detection performance of the proposed method, we compare its average run length (ARL) and expected detection delay (EDD) to the existing kernel-based methods presented in Table 1.444All code replicating our experiments is available in the supplement and at https://github.com/FlopsKa/rff-change-detection. For all experiments, we use the Gaussian kernel $K(\mathbf{x},\mathbf{y})=e^{-\gamma\left\|{\mathbf{x}-\mathbf{y}}\right\|_{2}^{2}}$ with $\gamma$ set by the median heuristic [12] or its RFF approximation, depending on the algorithm. All results were obtained on a PC with Ubuntu 20.04 LTS, 124GB RAM, and 32 cores with 2GHz each.
@@ 6.1 Synthetic data
Runtime. Figure 2 summarizes the runtime results of Algorithm 1 with the number of random Fourier features $r\in\{10,50,100,500,1\,000\}$ and for streams of length up to n=250 000n=250\,000. The experiments verify the $\mathcal{O}(r\log n)$ runtime complexity of the proposed algorithm, derived analytically in Section 4.2. We note that the dependence on $d$ is linear; we consider $d=1$ only.
@@
ARL vs. expected detection delay. To illustrate the EDD for a given target ARL, we reproduce the experiments of Wei and Xie [59, Figure 4], also taking our method into account. Specifically, we consider the pre-change distribution $\mathbb{P}=\mathcal{N}\!\left(\bm{0}_{20},\mathbf{I}_{20}\right)$ and set the parameters of each algorithm as follows. Matching the settings of the reproduced experiment, we choose $B_{\text{max}}=50$ and $N=15$ for online kernel CUSUM; for Scan B-statistics and NewMA, we set $B_{0}=50$. The remaining parameters of NewMA then follow from the heuristics detailed by the authors [24]. For Online RFF-MMD, we set r=1 000r=1\,000. We compute the thresholds for a given target ARL by processing $10\times(\text{target ARL})$ samples with each algorithm, repeating for $100$ Monte Carlo (MC) iterations, and computing the 11/(target ARL)1-1/($\text${target ARL}) quantile of the resulting test statistics. For approximating the EDD of each algorithm, we draw $64$ samples from $\mathbb{P}$, respectively, before sampling from $\mathbb{Q}$; we report the average over $100$ repetitions. OKCUSUM and ScanB additionally receive 1 0001\,000 samples from $\mathbb{P}$ upfront, to use as a reference sample. For NewMA, we process $400$ additional samples from $\mathbb{P}$ for both the MC estimate and the EDD experiment, to reduce its variance.
@@ 6.2 MNIST data
In this section, we interpret the MNIST data set [9] as high-dimensional data stream, similar to Wei and Xie [59, Figure 7], with the goal of detecting a change when the digit changes from 0 to a different digit. The experimental setup matches that of Section 6.1, but with $d=784$. Figure 4 collects our results; the results for the digits 4–6 are similar and in Appendix A.3.1. Similar to Figure 3, the proposed Online RFF-MMD algorithm shows very good performance. In particular, it achieves a lower EDD than all tested competitors throughout.
Figure 4: Average detection delay from MNIST digit 0 to digits 11, 22, and 33 (left to right).
@@ 7 Limitations
As with all kernel-based tests, the choice of kernel impacts the power of the test. While our theoretical guarantees (Section 4.3) hold for any kernel satisfying Assumption 1 and do not require any pre-change data, one usually selects the kernel or its parameters using a few available samples in practice. While kernel optimization is not the focus of this work, there exist works [20, 21, 22, 34, 50, 51, 16, 17, 18] to (approximately) achieve this goal; it is interesting future work to tackle this problem in the sequential setting. A separate future direction is considering non-i.i.d. data, for example, data that is strongly mixing [6] or data exhibiting functional dependence [60].
@@
The authors thank Zoltán Szabó and Tengyao Wang for helpful discussions. FK thanks Georg Gntuni and Marius Bohnert for exchanges on the algorithm’s implementation. This work was supported by the pilot program Core-Informatics of the Helmholtz Association (HGF).
@@ Appendix A Appendix
This appendix is structured as follows. We detail the extension of Online RFF-MMD to take a known or observed pre-change distribution into account in Appendix A.1. In particular, we show that in these settings tighter thresholds are possible. Appendix A.2 shows how to adapt our algorithm to detect multiple change points and gives corresponding guarantees. Additional numerical results of MMD-RFF are in Appendix A.3; we include numerical results w.r.t. our tighter thresholds in Appendix A.3.5. Appendix A.4 collects all our proofs, with auxiliary results in Appendix A.5 and external statements in Appendix A.6.
A.1 Known or estimable pre-change distribution
@@ A.1.1 Incorporating information of the pre-change distribution
To begin with, we consider the setting in which, for some $\nu\in\mathbb{N}$, historical data $X_{-\nu+1},\dots,X_{0}$ known to be from the (nonetheless unknown) pre-change distribution $\mathbb{P}$ is available. This setting has been studied in the literature from both the parametric [44] and non-parametric [59] perspectives. It is straightforward to extend the Online RFF-MMD stopping time defined in the main text to take advantage of the additional information. Intuitively, for each local test one may prepend the historical data to the block of data taken to be from the pre-change distribution. More formally, the following stopping time can be used:
$N=\inf\left\{n\geq 2\mid\bigcup_{j=0}^{\left\lfloor\log_{2}(n)\right\rfloor-1}\sqrt{\frac{2^{j}(n+\nu-2^{j})}{n+\nu}}\operatorname{MMD}_{\hat{K}}\!\left[X_{(-\nu+1):(n-2^{j})},X_{(n-2^{j}+1):n}\right]>\lambda_{n}\right\}.$
@@
$N=\inf\left\{n\geq 2\mid\bigcup_{j=0}^{\left\lfloor\log_{2}(n)\right\rfloor-1}\sqrt{2^{j}}\operatorname{MMD}_{\hat{K}}\!\left[\mathbb{P},X_{(n-2^{j}+1):n}\right]>\lambda_{n}\right\}$
@@ A.1.2 Sharper thresholds through knowledge of the pre-change distribution
The main tool for proving Theorems 1 and 2 is Lemma 5, which controls the tail behavior of the tests performed by Online RFF-MMD under their respective local nulls. With additional knowledge of the pre-change distribution, Lemma 5 can be significantly sharpened by taking the second moment of the feature map into account. To that end, we have the following result.
@@
$\mathbb{P}\!\left(\operatorname{MMD}_{\hat{K}}\!\left[X_{1:n},Y_{1:m}\right]>\varepsilon\right)\leq 4\exp\left(-\frac{1}{2}\min(m,n)\varepsilon^{2}\left[\tilde{\sigma}^{2}+2\varepsilon\right]^{-1}\right),$
where $\tilde{\sigma}^{2}=\mathbb{E}_{X\sim\mathbb{P}}\left[K\left(X,X\right)\right]-\mathbb{E}_{X,Y\sim\mathbb{P}}\left[K\left(X,Y\right)\right]$
@@
$\lambda_{n,j}$ $=\frac{2\sqrt{2}f^{2}(\gamma)}{\sqrt{\min(2^{j},n-2^{j})}}+\tilde{\sigma}f(\gamma),\hskip 20.00003ptn\in\mathbb{N},\;j\leq\log_{2}(n)$ (13)
$f(\gamma)$ $=\sqrt{2\log(16\gamma\log_{2}(2\gamma))}$
it holds that $\mathbb{E}_{\infty}[N]\geq\gamma$ where $N$ is as defined in (8).
@@
$\lambda_{n,j}=\frac{2\sqrt{2}f^{2}(\alpha,n)}{\sqrt{\min(2^{j},n-2^{j})}}+\tilde{\sigma}f(\alpha,n),\hskip 20.00003ptn\in\mathbb{N},j\leq\log_{2}(n)$
$f(\alpha,n)=\sqrt{2(\log(4n/\alpha)+2\log(\log_{2}(n))+\log(\log_{2}(2n)))}$
it holds that $\mathbb{P}_{\infty}(N<\infty)\leq\alpha$ where $N$ is as defined in (8).
@@
A.2 Online detection of multiple change points
In this section, we explain how the algorithm in the main part can be extended to detect multiple change points in an online fashion. Indeed, consider the setting where data $X_{1},X_{2},\dots$ are observed sequentially. There is a sequence of integers $(\eta_{k})_{k\in\mathbb{N}}$ with $\eta_{0}=0$ and $\eta_{k}<\eta_{k+1}$ for each $k\in\mathbb{N}$, and a sequence of measures $(\mathbb{Q}_{k})_{k\in\mathbb{N}}$ each in $\mathcal{M}^{1}_{+}$ with $\mathbb{Q}_{k}\neq\mathbb{Q}_{k+1}$ for each $k\in\mathbb{N}$ such that for each $t\in\mathbb{N}$ the data are distributed according to
$X_{t}\sim\left\{\begin{matrix}\mathbb{Q}_{1}&\text{if}\quad\eta_{0}<t\leq\eta_{1}\\ \mathbb{Q}_{2}&\text{if}\quad\eta_{1}<t\leq\eta_{2}\\ \vdots&\\ \end{matrix}\right.$.
Consider running Algorithm 1 on this data stream using the threshold sequence given in Theorem 2. As soon as a change is detected, say at location $n$, we re-start the algorithm at $X_{n+1}$. However, the index on the threshold sequence is maintained. More precisely: having detected a change at time $n$, at the $(n+j)$-th iteration (for $j\geq 1$), we construct local statistics over the interval $\{n+1,\dots,n+j\}$ according to Section 4.1, however the local statistics continue to be compared with the threshold $\lambda_{n+j}$. We are able to prove the following guarantee on the change point locations recovered in this way:
Theorem 5.
For any deterministic time $T<\infty$ let $M$ denote the number of $\eta_{k}$-s taking values in $\left\{1,\dots,T\right\}$. Write $\delta_{k}=\min\left(\eta_{k}-\eta_{k-1},\eta_{k+1}-\eta_{k}\right)$ for $k=1,\dots,M-1$ and $\delta_{M}=\min\left(\eta_{M}-\eta_{M-1},T-\eta_{M}\right)$ for the effective sample size associated with the $k$-th change point location. For some $\alpha\in(0,1)$, let $\hat{M}$ be the number of change points detected up to time $T$ using the threshold sequence given in Theorem 2 and the procedure described above, and let $\hat{\eta}_{1},\dots,\hat{\eta}_{\hat{M}}$ be their locations. If $\cup_{k=1}^{M}\operatorname{supp}\left(\mathbb{Q}_{k}\right)\subseteq\mathcal{X}$ for some compact set $\mathcal{X}\subset\mathbb{R}^{d}$,
$\delta_{k}\geq C_{1}\frac{\log\left(2T/\alpha\right)}{\left(\operatorname{MMD}_{K}\left[\mathbb{Q}_{k},\mathbb{Q}_{k+1}\right]\right)^{2}}\quad\text{for}\quad k=1,\dots,M,$
and the number of random Fourier features is chosen so that
$\sqrt{r}\geq C_{2}\frac{h\!\left(d,\left|\mathcal{X}\right|,\sigma\right)+\sqrt{2\log\left(2/\alpha\right)}}{\left(\min_{k=1,\dots M}\operatorname{MMD}_{K}\!\left[\mathbb{Q}_{k},\mathbb{Q}_{k+1}\right]\right)^{2}},$
then with probability at least $1-2\alpha$ it holds that $\hat{M}=M$ and
$\eta_{k}<\hat{\eta}_{k}\leq\eta_{k}+C_{3}\frac{\log\left(2T/\alpha\right)}{\left(\operatorname{MMD}_{K}\left[\mathbb{Q}_{k},\mathbb{Q}_{k+1}\right]\right)^{2}}\quad\text{for}\quad k=1,\dots M.$
Here $h\!\left(d,\left|\mathcal{X}\right|,\sigma\right)$ is as given in Theorem 3, and $C_{1}$, $C_{2}$, and $C_{3}$ are absolute constants independent of $\alpha$ and the sequences $\{\delta_{k}\}_{k=1,\dots,M}$ and $\{\operatorname{MMD}_{K}\!\left[\mathbb{Q}_{k},\mathbb{Q}_{k+1}\right]\}_{k=1,\dots,M}$.
@@ A.3 Additional experiments
In this section, we summarize additional numerical results. The MNIST results with the MC threshold as in the main text are in Section A.3.1. We include results obtained without threshold estimation in Section A.3.4. A comparison of the distribution dependent and distribution-free bounds is in Section A.3.5. Numerical studies on the Human Activity Sensing Consortium (HASC) and on the MarzukaBL data sets are in Section A.3.2 and in Section A.3.3, respectively.
A.3.1 MNIST data digits 4–9
In the following Figure 5, we collect the change detection performances of the algorithms of Table 1 on MNIST data, where the digit changes from 0 to one of 4499; the results on the other digits and the experimental setup are in Section 6.2. As in the corresponding experiment in the main part of this article, our proposed Online RFF-MMD algorithm consistently achieves the lowest expected detection delay, highlighting its good practical performance.
@@ A.3.1 MNIST data digits 4–9
A.3.2 HASC (Human Activity Sensing Consortium) data
This section collects our experiments on the Human Activity Sensing Consortium (HASC; available at http://hasc.jp/hc2011/) challenge $2011$ data set, which is also considered in Liu et al. [35], Li et al. [33], Wei and Xie [59]. As in Wei and Xie [59], we consider the change from walking to staying of participant $101$. For preprocessing, we order the corresponding csv-files in the data set lexicographically, omitting the first 1 5961\,596 (detailed below) samples of walking, and then concatenating $100$ walking observations and $100$ staying observations to obtain a total of $10$ data sets (with $d=3$) with a single change point each. For obtaining the thresholds for the proposed Online RFF MMD, NewMA, Scan-B statistics, and OKCUSUM change detection algorithms, we proceed as elaborated in Section 6.1, that is, we use a Monte Carlo approach with $\alpha=0.1$ on all $10$ walking data sets. As per Wei and Xie [59], for ScanB and OKCUSUM, we set the number of windows $N=14$ and the window length $w=114$, implying that both algorithms receive $14\cdot 114=1\,596$ samples from walking upfront. Likewise, we process 1 5961\,596 elements with NewMA upfront. All kernel-based approaches use the (approximated) Gaussian kernel with the $\gamma>0$ parameter set by the median heuristic.
For the density ratio-based (i.e., non-kernel-based) RuLSIF algorithm—which showed the best performance on HASC in Liu et al. [35]—, we use the python changepoynt implementation and consider the $l_{2}$-norm of each three-dimensional observation. [^5] The changepoynt library is available at https://github.com/Lucew/changepoynt. We then obtain the change scores for the full data set and select the point with the highest score as predicted change point, as done in Liu et al. [35, Section 4.2]. To match their setup, we set the window length to $50$, the number of windows to $10$, and $\alpha=0.1$. Additionally, due to the large number of “too early” cases reported in this setup, we introduce an offset and take the window length to be equal to the offset. In this setup, we consider the point with the highest score plus the offset as reported change point. A “miss” can not occur, due to the scores reported to RuLSIF having at least one maximum. Table 2 collects our results.
Table 2: Comparison of change detection algorithms on $10$ data sets derived from the HASC data set. “Too early” refers to the number of times an algorithm reported a change point before the actual change occurred. “Miss” reports cases in which no change was reported.
Algorithm Average delay Too early Miss
Online RFF MMD 21.86 2 1
NewMA 34.25 1 5
ScanB 31.20 0 0
OKCUSUM 17.44 1 0
RuLSIF 4.50 8 0
RuLSIF ($\text{offset}=30$) 20.38 2 0
RuLSIF ($\text{offset}=40$) 35.0 3 0
RuLSIF ($\text{offset}=50$) 39.2 0 0
We emphasize that in the above case all algorithms except for the proposed one and RuLSIF receive 1 5961\,596 samples upfront, to use as a reference sample. Even though this fundamentally favors the kernel-based competitors, our proposed method achieves results that are comparable to those of the other kernel-based approaches and to RuLSIF.
A.3.3 MazurkaBL data
In Kosta et al. [26], the authors found that change points in loudness information of Chopin’s Mazurkas correspond to score positions having dynamic markings, tempo, or expression markings, among others. To further validate our algorithm’s performance, we additionally run the proposed method on the “M17-4” sample (illustrated in Figure 6) of pianist “pid50534-05” of their MazurkaBL [27] data set with the goal of detecting these changes. Here, the dimensionality is $d=1$.
Figure 6: An excerpt of Frédéric Chopin’s Mazurka Op. 17 No. 4.
As in the main article (see Section 6), we approximate the Gaussian kernel with $\gamma>0$ set by the median heuristic. For obtaining the thresholds using Monte Carlo iterations, we slice the data along each annotated change point, where we consider the annotations provided by Kosta et al. [27] as ground truth, and compute the test statistics obtained by the proposed Online RFF MMD algorithm individually on each one. We then select the 10.11-0.1-quantile across all test statistics so obtained as threshold.
To detect change points, we again slice the loudness data, but now such that each slice contains precisely one change point, which yields a total of $25$ samples with an average length of 1 561.321\,561.32. We process each one of these with our proposed method and consider the first time the test statistic exceeds the threshold as “change”. In total, our proposed method flags $10$ change points too early, and, on the remaining $15$ has an average detection delay of $73.67$, with a median detection delay of $64.0$.
While, when contrasting these results with the results obtained on the MNIST (Section A.3) and HASC (Section A.3.2) data sets, detecting changes in the selected loudness data of a Mazurka seems substantially more challenging, the proposed algorithm still manages to detect many changes with a relatively small delay.
A.3.4 Distribution-free bound